Circumference of circle of radius 2.1cm using significant figures
1
Expert's answer
2020-04-23T17:41:59-0400

The circumference of circle is "2\pi r", where r is the radius of the circle.

Given, r = 2.1cm.

Circumference of circle "= 2 * \pi * 2.1 = 2 * 3.1416* 2.1= 13.1947"